# On the reconstruction map in JT gravity

### v1

@@ 4.3 Numerical simulations
For numerical simulations, we generate a spectrum with the JT density of states given by (1.5) and CUE level correlations, by using (3.28) and (3.30) to “fold” a CUE spectrum (which has a uniform density of states on $[0,2\pi)$) to have the JT density of states. First, we define a smooth folding function $\overline{F}$ for an “averaged” spectrum via a variant of (3.30):
@@
Due to numerical limitations, our simulations necessarily truncate the spectrum to a finite number of levels $D$($=5000$). For the $x$ operator, we use a close approximation to (3.36):
@@
Finally, for the initial state itself, we will illustrate two choices: the Hartle-Hawking state in (1.6) with a specific value of $\beta$, truncated to $D$ levels, as well as a microcanonical state in the interval $[E(\beta),3E(\beta)]$ supported over $d$ energy levels near the peak of the Hartle-Hawking distribution. Consequently, both states should show similar behavior in the “most probable” regions of their respective distributions, but the Hartle-Hawking state should show a larger spread in velocities and positions. Here, it is important to emphasize a limitation of our finite size numerics: due to the small number of energy levels $D$ and the exponentially growing density of states, we cannot simultaneously ensure that $d$ is large for the microcanonical state and that $\rho(E)$ remains approximately constant (i.e., satisfies (4.4)) over its support. We have prioritized the former, and our “microcanonical state” is supported on $d=595$ energy levels, but $\rho(3E(\beta))$ is several times larger than $\rho(E(\beta))$. This is still a much narrower spread than the Hartle-Hawking state (in particular, the microcanonical state lacks a long tail), and our numerical results still effectively illustrate the qualitative differences in wormhole dynamics between the two initial states that originate in the significantly larger energy spread of the Hartle-Hawking state.
@@
In contrast to [11], our reconstruction map suggests that the wormhole length is sharply defined all the way up to $t\sim\mathrm{e}^{S}/\sqrt{S}$. Our result is a consequence of our guiding principle that semiclassical dynamics should be as accurate as possible at short times.
@@ 5 Velocity operator
First, notice that so long as $x$ is a real Hermitian operator, $v$ is antisymmetric and imaginary (thus Hermitian). As $v$ is antisymmetric, any eigenvalue $\lambda$ has a partner eigenvalue $-\lambda$; since it is Hermitian, these eigenvalues are real. We deduce that when measuring the velocity operator $v$ in a typical state, as one would expect to find at times $t\gg\mathrm{e}^{S}$, we have a 50% chance of measuring $v<0$ and a 50% chance of measuring $v>0$.
One can wonder whether a stronger statement can be made. Perhaps the probability to measure $v\approx\sqrt{2E}$ is itself about 50%, as is the probability to measure $v\approx-\sqrt{2E}$ (see e.g. discussion in [1]). A heauristic argument for this is that a general state of approximate energy $E$ can be written as a superposition of Hartle-Hawking states evolved for times approximately within $t\in[-\mathrm{e}^{S(E)},\mathrm{e}^{S(E)}]$. Semiclassically these states are all either shrinking with velocity $-\sqrt{2E}$ or growing with velocity $\sqrt{2E}$. The question is: how well does the reconstructed velocity operator match this semiclassical answer? This is a quantitative prediction that we can address. For simplicity, let us focus on the microcanonical state discussed in Section 4.1.
@@
These results imply that the probability distribution of $v$ at a typical late time has an O(1) spread around ±$\pm\sqrt{2E}$. We deduce that the distribution of $v^{2}$ cannot be too sharply peaked at $2E$.
@@ 6 Outlook
We have proposed an explicit reconstruction map for pure JT gravity, with the bulk/effective theory given by JT gravity at $S_{0}\to\infty$ [5] and the boundary/fundamental theory given by a single draw of the matrix model dual [7]. As a case study, we wrote down an explicit proposal for the wormhole length operator in the dual theory and studied its behavior as a function of time in (initially semiclassical) states of interest in quantum gravity.
@@ 6.1 How unique are the predictions of our reconstruction map?
for some function $g(\vartheta)=\vartheta+\cdots$ whose Taylor expansion matches the semiclassical expectations at short times, but differs at long times. The arguments of Appendix A generalize to this case, and we still find that $\langle x_{\mathrm{alt}}(t)\rangle$ (using obvious abbreviated notation) overshoots its late time value, and that the wave function exhibits an explosive growth in fluctuations just before the Heisenberg time. An explicit example replaces (3.36) with a formula much closer to [9]:
@@
We have determined these matrix elements according to certain consistency conditions. The off-diagonal elements for large $(n-m)$ must have the same average (over some range of $(n-m)$) as those of $R^{*}(x)$ to reproduce the correct semiclassical dynamics of the velocity $v(t)$, which is responsible for the $(-1)$ prefactor above. The diagonal elements are chosen to ensure that $x_{\mathrm{alt}}(0)\approx 0$ in e.g. the microcanonical or Hartle-Hawking states. Moreover, to satisfy the guiding principle of Sec. 3, the off-diagonal matrix elements reduce to functions of $(n-m)$ in microcanonical windows, rather than e.g. $(E_{n}-E_{m})$ as in [1, 9]. [^7] This discrepancy is rather small due to the spectral rigidity, once $|m-n|\gg 1$.
The alternate length operator $x_{\mathrm{alt}}(t)$ exhibits extremely similar dynamics to that observed in Figure 1: see Figure 4. This happens because $R^{*}_{\mathrm{alt}}(x)$ (in microcanonical windows) is still an even function of $(n-m)$; it simply arises from a different choice of $g(\vartheta)$ in (6.1).
@@ 6.2 Complexity = volume?
The key outcome of our construction is that the reconstructed length operator could be studied in detail using the theory of quantum ergodicity [4]. This allowed us to leverage many known results from random matrix theory to make clear predictions for the dynamics of wormholes before and after the Heisenberg time. The most striking of these predictions are a non-monotonic growth in the averaged wormhole length operator, along with heavy-tailed moments of the length operator in the Hartle-Hawking states. As we have emphasized above, we believe that this prediction is generic and is not sensitive to particular assumptions in our reconstruction procedure.
@@
What are the implications for the “complexity = volume” conjecture [8]? Often “complexity” is assumed to mean circuit complexity, the minimal number of gates neeed to prepare the state from some reference state. It is unclear that this conjecture for circuit complexity is consistent with the non-monotonic growth.
@@
That said, it is very intriguing to compare our result to an independent result which has recently appeared in the literature [14] regarding a different notion of complexity. These authors pointed out that the “spread complexity” [21] of a thermofield double state exhibits a similar non-monotonic dependence in time as our $\langle x(t)\rangle$ – we believe that the origin of this phenomena is analogous to what we have shown: the wormhole shrinks at $t\sim t_{\mathrm{H}}$ because of level repulsion in the spectrum of the fundamental theory. The authors of [14] have proposed that spread complexity describes the length operator in quantum gravity. We have arguably “derived” a closely related equivalence between wormhole length and quantum ergodicity in our formalism. A clear advantage, in our view, of our approach to defining the length operator is that the spread complexity is based on a notion of Krylov complexity [22, 23] which is sensitive to initial states; our approach defines the length operator independently of the initial state.
@@
Interestingly, these results have implications for the non-gravitational physics of quantum many-body systems. We have shown, with very reasonable assumptions, that ergodic dynamics in the fundamental (matrix) theory drives the dynamics of observables in JT gravity such as the wormhole length over the Heisenberg timescale $t_{{\rm H}}\sim\exp(S(E))$. In turn, the “complexity = volume” conjecture [8] posits a relation between the length operator, together with a possibly larger class of gravitational observables with nontrivial dynamics over this timescale [24], and the growth of the (e.g., spread) complexity of a state in the fundamental theory over the same timescale. This suggests that ergodicity may be the physical mechanism driving the dynamics of non-gravitational observables such as complexity in the fundamental theory over the Heisenberg timescale, which may be of relevance in suitable classes of quantum systems.
@@ 6.3 The linearity of $R^{*}$ and adding matter
A peculiar feature of our reconstruction map is that it is linear. It is well-known that a linear reconstruction map $R^{*}$ cannot be good if $V$ is non-isometric, i.e. it cannot satisfy $R^{*}(O)V\ket{\psi}\simeq VO\ket{\psi}$ to a good approximation for all |$\ket{\psi}$ and even just simple $O$ [13, 25]. Here we evade the issue because we do not demand the reconstruction is good on all states. We only demand it on states that we call “semiclassical”. For example, we do not require it works on highly entangled states like $\int\mathrm{d}x\,\ket{x}_{0}\ket{x}_{R}$, where $R$ is some auxiliary reference system and perhaps the integral runs over some finite range of $x$. This state is intuitively non-semiclassical, because $R$ has decohered the wormhole length into a incoherent mixture of very high (non-semiclassical) precision. It is these kinds of highly entangled states that easily lead to issues with linear reconstruction maps in non-isometric codes [13, 25].
This situation will be harder once matter is included [1, 26, 27, 28]. Then it will be possible to highly entangle the matter with a reference system while keeping the state “semiclassical”. For this reason, once matter is included, we expect a linear reconstruction map will have to be abandoned.
@@ Appendix A Quantum dynamical ergodicity
This appendix contains a review of technical details of the theory of quantum dynamical ergodicity [4], which play a crucial role in our choice of reconstruction $R^{*}$ and ultimate predictions for wormhole dynamics.
@@ A.1 Microcanonical wavepackets and quantum ergodicity
Here we provide a more formal discussion of quantum ergodicity [4] within microcanonical windows, and justify how our reconstructing map satisfies the guiding principle in Sec. 3. Because the $E_{n}$ have an irregular and bounded spectrum, we cannot expect that any Hermitian operator $\delta$ generates exact time-translations, as emphasized in the main text. Such a conjugate relationship requires precise action-angle variables, such as $(N,\theta)$ defined in the main text, whose analogues in the classical phase space are given by the expressions (from (3.26) and (3.27)):
$N$ $=\int_{0}^{H}\mathrm{d}E\ \rho(E)~{},$ (A.1a)
$\theta$ $=\frac{\delta}{\rho(H)}~{},$ (A.1b)
where $\rho(E)$ is the (averaged) density of states.
Still, we might hope for $(H,\delta)$ to effectively function as a good action-angle pair at short times in sufficiently narrow energy windows. More explicitly, consider some narrow range of energies $[E_{\text{mc}},E_{\text{mc}}+\Delta E]$ such that the averaged density of states is constant,
@@
$N$ $\approx H\rho(E_{\text{mc}})+\text{const.}~{},$. , (A.3a)
$\theta$ $\approx\frac{\delta}{\rho(E_{\text{mc}})}~{}.$ (A.3b)
@@
Figure 5: Probability disributions $\lvert\langle\theta^{\prime}|\tilde{\theta}\rangle\rvert^{2}$ for the wavepackets in (A.8) for different values of $-\pi\leq\theta\leq\pi$ in $\theta^{\prime}\in(-\pi,\pi]$, with $d=60$. These wavepackets become increasingly localized as $d\to\infty$, and always contain a complete orthonormal basis for any finite $d$. The $N_{0}$-dependent oscillations in the phase of the wavepackets are not visible in this plot of probability density, but will be visible in the real and imaginary parts.
@@
Imposing (A.10) for small $\theta_{0}$ implies [4] the following constraints:
@@
    $N$ is an exactly conserved quantity within the microcanonical window:
    $[N_{d},H_{d}]=0,$ (A.11)
@@
For other (suboptimal) choices $(N^{\prime},\theta^{\prime})$, the long-time dynamics of $\delta$ would generally be drastically different even to leading order. However, any global reconstruction of the above form continues to be sensitive to the optimal microcanonical dynamics of $\theta$ identified by our physical principle at all times, to within some resolution $\Delta\theta\ll 1$ determined by the approximation in (A.14).

### v2

@@ 4.3 Numerical simulations
For numerical simulations, we generate a spectrum with the JT density of states given by (1.5) and CUE level correlations, by using (3.28) and (3.30) to “fold” a CUE spectrum (which has a uniform density of states on $0,2\pi)$) to have the JT density of states. First, we define a smooth folding function $\overline{F}$ for an “averaged” spectrum via a variant of (3.30):
@@
Due to numerical limitations, our simulations necessarily truncate the spectrum to a finite number of levels DD(=5000=5000). For the $x$ operator, we use a close approximation to (3.36):
@@
Finally, for the initial state itself, we will illustrate two choices: the Hartle-Hawking state in (1.6) with a specific value of $\beta$, truncated to $D$ levels, as well as a microcanonical state in the interval $E(\beta),3E(\beta)]$ supported over $d$ energy levels near the peak of the Hartle-Hawking distribution. Consequently, both states should show similar behavior in the “most probable” regions of their respective distributions, but the Hartle-Hawking state should show a larger spread in velocities and positions. Here, it is important to emphasize a limitation of our finite size numerics: due to the small number of energy levels $D$ and the exponentially growing density of states, we cannot simultaneously ensure that $d$ is large for the microcanonical state and that $\rho(E)$ remains approximately constant (i.e., satisfies (4.4)) over its support. We have prioritized the former, and our “microcanonical state” is supported on $d=595$ energy levels, but $\rho(3E(\beta))$ is several times larger than $\rho(E(\beta))$. This is still a much narrower spread than the Hartle-Hawking state (in particular, the microcanonical state lacks a long tail), and our numerical results still effectively illustrate the qualitative differences in wormhole dynamics between the two initial states that originate in the significantly larger energy spread of the Hartle-Hawking state.
@@
In contrast to [Miyaji:2024ity], our reconstruction map suggests that the wormhole length is sharply defined all the way up to $t\sim\mathrm{e}^{S}/\sqrt{S}$. Our result is a consequence of our guiding principle that semiclassical dynamics should be as accurate as possible at short times.
@@ 5 Velocity operator
First, notice that so long as $x$ is a real Hermitian operator, $v$ is antisymmetric and imaginary (thus Hermitian). As $v$ is antisymmetric, any eigenvalue $\lambda$ has a partner eigenvalue $\lambda$; since it is Hermitian, these eigenvalues are real. We deduce that when measuring the velocity operator $v$ in a typical state, as one would expect to find at times $t\gg\mathrm{e}^{S}$, we have a 50% chance of measuring $v<0$ and a 50% chance of measuring $v>0$.
One can wonder whether a stronger statement can be made. Perhaps the probability to measure $v\approx\sqrt{2E}$ is itself about 50%, as is the probability to measure $v\approx-\sqrt{2E}$ (see e.g. discussion in [Iliesiu:2024cnh]). A heauristic argument for this is that a general state of approximate energy $E$ can be written as a superposition of Hartle-Hawking states evolved for times approximately within $t\in[-\mathrm{e}^{S(E)},\mathrm{e}^{S(E)}]$. Semiclassically these states are all either shrinking with velocity $\sqrt{2E}$ or growing with velocity $\sqrt{2E}$. The question is: how well does the reconstructed velocity operator match this semiclassical answer? This is a quantitative prediction that we can address. For simplicity, let us focus on the microcanonical state discussed in Section 4.1.
@@
These results imply that the probability distribution of $v$ at a typical late time has an O(1) spread around $\pm\sqrt{2E}$. We deduce that the distribution of $v^{2}$ cannot be too sharply peaked at $2E$.
@@ 6 Outlook
We have proposed an explicit reconstruction map for pure JT gravity, with the bulk/effective theory given by JT gravity at $S_{0}\to\infty$ [Harlow:2018tqv] and the boundary/fundamental theory given by a single draw of the matrix model dual [Saad:2019lba]. As a case study, we wrote down an explicit proposal for the wormhole length operator in the dual theory and studied its behavior as a function of time in (initially semiclassical) states of interest in quantum gravity.
@@ 6.1 How unique are the predictions of our reconstruction map?
for some function $g(\vartheta)=\vartheta+\cdots$ whose Taylor expansion matches the semiclassical expectations at short times, but differs at long times. The arguments of Appendix A generalize to this case, and we still find that $\langle x_{\mathrm{alt}}(t)\rangle$ (using obvious abbreviated notation) overshoots its late time value, and that the wave function exhibits an explosive growth in fluctuations just before the Heisenberg time. An explicit example replaces (3.36) with a formula much closer to [Iliesiu:2021ari]:
@@
We have determined these matrix elements according to certain consistency conditions. The off-diagonal elements for large (nm)(n-m) must have the same average (over some range of (nm)(n-m)) as those of $R^{*}(x)$ to reproduce the correct semiclassical dynamics of the velocity $v(t)$, which is responsible for the (1)(-1) prefactor above. The diagonal elements are chosen to ensure that $x_{\mathrm{alt}}(0)\approx 0$ in e.g. the microcanonical or Hartle-Hawking states. Moreover, to satisfy the guiding principle of Sec. 3, the off-diagonal matrix elements reduce to functions of (nm)(n-m) in microcanonical windows, rather than e.g. $(E_{n}-E_{m})$ as in [Iliesiu:2024cnh, Iliesiu:2021ari]. [^7] This discrepancy is rather small due to the spectral rigidity, once $|m-n|\gg 1$.
The alternate length operator $x_{\mathrm{alt}}(t)$ exhibits extremely similar dynamics to that observed in Figure 1: see Figure 4. This happens because $R^{*}_{\mathrm{alt}}(x)$ (in microcanonical windows) is still an even function of (nm)(n-m); it simply arises from a different choice of $g(\vartheta)$ in (6.1).
In this context, we also comment on an interesting open problem. The effective theory, due to the $S_{0}\to\infty$ limit, corresponds entirely to the leading cylinder (single wormhole) topology in the gravitational path integral. However, the fundamental theory involves a sum over topologies, but our construction only identifies abstract operators acting on the fundamental Hilbert space without a direct geometric interpretation. In other words, we have argued that $R^{*}(x)$ corresponds to the wormhole length in the leading topology at short times, but it would be nice to have a more explicit understanding of which geometric quantities it corresponds to in the full nonperturbative path integral. A full understanding should resolve the ambiguities highlighted here.
@@ 6.2 Complexity = volume?
The key outcome of our construction is that the reconstructed length operator could be studied in detail using the theory of quantum ergodicity [dynamicalqergodicity]. This allowed us to leverage many known results from random matrix theory to make clear predictions for the dynamics of wormholes before and after the Heisenberg time. The most striking of these predictions are a non-monotonic growth in the averaged wormhole length operator, along with heavy-tailed moments of the length operator in the Hartle-Hawking states. As we have emphasized above, we believe that this prediction is generic and is not sensitive to particular assumptions in our reconstruction procedure.
@@
What are the implications for the “complexity = volume” conjecture [Susskind:2014rva]? Often “complexity” is assumed to mean circuit complexity, the minimal number of gates needed to prepare the state from some reference state. It is unclear that this conjecture for circuit complexity is consistent with the non-monotonic growth.
@@
That said, it is very intriguing to compare our result to an independent result which has recently appeared in the literature [Balasubramanian:2024lqk] regarding a different notion of complexity. These authors pointed out that the “spread complexity” [spreadcomplexity] of a thermofield double state exhibits a similar non-monotonic dependence in time as our $\langle x(t)\rangle$ – we believe that the origin of this phenomena is analogous to what we have shown: the wormhole shrinks at $t\sim t_{\mathrm{H}}$ because of level repulsion in the spectrum of the fundamental theory. The authors of [Balasubramanian:2024lqk] have proposed that spread complexity describes the length operator in quantum gravity; see also [SonnerDSSYK]. We have arguably “derived” a closely related equivalence between wormhole length and quantum ergodicity in our formalism. A clear advantage, in our view, of our approach to defining the length operator is that the spread complexity is based on a notion of Krylov complexity [Parker:2018yvk, KrylovReview] which is sensitive to initial states; our approach defines the length operator independently of the initial state.
@@
Interestingly, these results have implications for the non-gravitational physics of quantum many-body systems. We have shown, with very reasonable assumptions, that ergodic dynamics in the fundamental (matrix) theory drives the dynamics of observables in JT gravity such as the wormhole length over the Heisenberg timescale $t_{{\rm H}}\sim\exp(S(E))$. In turn, the “complexity = volume” conjecture [Susskind:2014rva] posits a relation between the length operator, together with a possibly larger class of gravitational observables with nontrivial dynamics over this timescale [ComplexityAnything], and the growth of the (e.g., spread) complexity of a state in the fundamental theory over the same timescale. This suggests that ergodicity may be the physical mechanism driving the dynamics of non-gravitational observables such as complexity in the fundamental theory over the Heisenberg timescale, which may be of relevance in suitable classes of quantum systems.
@@ 6.3 The linearity of $R^{*}$ and adding matter
[…]
This situation will be harder once matter is included [Iliesiu:2024cnh, Penington:2019kki, Penington:2023dql, Boruch:2024kvv]. Then it will be possible to highly entangle the matter with a reference system while keeping the state “semiclassical”. For this reason, once matter is included, we expect a linear reconstruction map will have to be abandoned.
@@ Appendix A Quantum dynamical ergodicity
This appendix contains a review of technical details of the theory of quantum dynamical ergodicity [dynamicalqergodicity], which play a crucial role in our choice of reconstruction $R^{*}$ and ultimate predictions for wormhole dynamics.
@@ A.1 Microcanonical wavepackets and quantum ergodicity
Here we provide a more formal discussion of quantum ergodicity [dynamicalqergodicity] within microcanonical windows, and justify how our reconstructing map satisfies the guiding principle in Sec. 3. At an intuitive level, the goal of this subsection is to associate the time-shift operator $\delta$ with the basis of discrete Fourier transforms of the energy eigenstates $\lvert E_{n}\rangle$ in narrow energy windows, owing to the expected conjugate relationship between $(E,\delta)$ that is represented in standard quantum mechanics by Fourier transforms. This will end up implying that the Hamiltonian $H$ acts as close as possible to a translation of $\delta$, and in this way satisfies our guiding principle.
Such an association may be formulated exactly for perfectly regular spectra such as in the harmonic oscillator. But in the case of JT gravity, because the $E_{n}$ have an irregular and bounded spectrum, we cannot expect that any Hermitian operator $\delta$ generates exact time-translations, as emphasized in the main text. Such a conjugate relationship requires precise action-angle variables, such as $(N,\theta)$ defined in the main text, whose analogues in the classical phase space are given by the expressions (from (3.26) and (3.27)):
${N}_{\text{cl}}$ $=\int_{0}^{{H}_{\text{cl}}}\mathrm{d}E\ \rho(E)~,$ (A.1a)
${\theta}_{\text{cl}}$ $=\frac{{\delta}_{\text{cl}}}{\rho(H)}~,$ (A.1b)
where $\rho(E)$ is the (averaged) density of states. Here, we use the notation ${A}_{\text{cl}}$ to indicate that a variable $A$ is classical, rather than in the (quantum) effective or fundamental theories.
Still, we might hope for $(H,\delta)$ to effectively function as a good action-angle pair at short times in sufficiently narrow energy windows. More explicitly, consider some narrow range of energies $E_{\text{mc}},E_{\text{mc}}+\Delta E]$ such that the averaged density of states is constant,
@@
$N$ $\approx H\rho(E_{\text{mc}})+\text{const.}~,$ (A.3a)
$\theta$ $\approx\frac{1}{\rho(E_{\text{mc}})}\delta~.$ (A.3b)
@@
Figure 5: Probability disributions $\lvert\langle\theta^{\prime}|\tilde{\theta}\rangle\rvert^{2}$ for the wavepackets in (A.8) for different values of $\pi\leq\theta\leq\pi$ in $\theta^{\prime}\in(-\pi,\pi]$, with $d=60$. These wavepackets become increasingly localized as $d\to\infty$, and always contain a complete orthonormal basis for any finite $d$. The $N_{0}$-dependent oscillations in the phase of the wavepackets are not visible in this plot of probability density, but will be visible in the real and imaginary parts.
@@
Imposing (A.10) for small $\theta_{0}$ implies [dynamicalqergodicity] the following constraints:
@@
    NN is an exactly conserved quantity within the microcanonical window:
    $N_{d},H_{d}]=0,$ (A.11)
@@
For other (suboptimal) choices $(N^{\prime},\theta^{\prime})$, the long-time dynamics of $\delta$ would generally be drastically different even to leading order. However, any global reconstruction of the above form continues to be sensitive to the optimal microcanonical dynamics of $\theta$ identified by our physical principle at all times, to within some resolution $\Delta\theta\ll 1$ determined by the approximation in (A.14). We note that in an entirely analogous manner, our guiding principle ensures that the length operator essentially reduces to the absolute value of the angle variable $\lvert\theta\rvert$ in microcanonical windows, via the expression in (3.34).